4. Suppose the production function is equal to the following: Q = (L)(K) Suppose the price of capital is equal to r, the price of labor is equal to w, and capital is fixed at 10 units.

a) Determine the Cost function.

b) Determine the marginal cost of producing an additional unit of output.

c) Determine the average variable cost.
Answer #1

Q = Tk Q=r (10) r = Q lo L-Q2 100 Cost = wh tok = wa² + lor loo b) marginal cost = 260 woo c) Average variable cost = wl - wo